OpenStudy (anonymous):

y varies directly as the square of x. When x = 2, y = 20. Find y when x = 4. y = 40 y = 80 y = 100 y = 120

OpenStudy (anonymous):

y varies directly as the square of x. When x = 2, y = 20. Find y when x = 4. Y= Cx^2 20=C2^2 C=5 y=5x^2 y= 5(4^2)= 5*16 =80